OpenStudy (anonymous):

State whether the function is odd, even, or neither. Solve algebraically. y=x/e^x

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

it is neither

OpenStudy (anonymous):

An even function is a function, f, for which f ( - x) = f (x). The graph of an even function is symmetric about the y-axis. An odd function is a function, f, with the property that f(-x)=-f(x). So look at the graph of y=x/e^x, and you tell me.

OpenStudy (anonymous):

.. Im still lost.

OpenStudy (anonymous):

Plug in any x in for the equation, and tell me if f(x) = f(-x), or if f(-x) = - f (x), or neither. If you do not understand this notation, or something else, tell me!

OpenStudy (anonymous):

i just dont understand the whole concept of e in the equation.

OpenStudy (anonymous):

e is just a constant, like pi. e is 2.71828183. . . Don't let it confuse you. The main point of the question at hand is just figuring out if the equation is even or odd. The same principals apply. If even, f(x) = f(-x) If odd, f(-x) = -f(x)

OpenStudy (anonymous):

Plug in any number. 3 for example. Is f(3) equal to f(-3)? 3/e^3 = .150. . . -3/e^-3 = -60.257. . . So, the function is not even. Is it odd? Or neither?

OpenStudy (anonymous):

neither! gracias.

OpenStudy (anonymous):

^_^ You're welcome.